### v1

# FAST-Splat: Fast, Ambiguity-Free Semantics Transfer in Gaussian Splatting

Abstract
We present FAST-Splat for fast, ambiguity-free semantic Gaussian Splatting, which seeks to address the main limitations of existing semantic Gaussian Splatting methods, namely: slow training and rendering speeds; high memory usage; and ambiguous semantic object localization. In deriving FAST-Splat, we formulate open-vocabulary semantic Gaussian Splatting as the problem of extending closed-set semantic distillation to the open-set (open-vocabulary) setting, enabling FAST-Splat to provide precise semantic object localization results, even when prompted with ambiguous user-provided natural-language queries. Further, by exploiting the explicit form of the Gaussian Splatting scene representation to the fullest extent, FAST-Splat retains the remarkable training and rendering speeds of Gaussian Splatting. Specifically, while existing semantic Gaussian Splatting methods distill semantics into a separate neural field or utilize neural models for dimensionality reduction, FAST-Splat directly augments each Gaussian with specific semantic codes, preserving the training, rendering, and memory-usage advantages of Gaussian Splatting over neural field methods. These Gaussian-specific semantic codes, together with a hash-table, enable semantic similarity to be measured with open-vocabulary user prompts and further enable FAST-Splat to respond with unambiguous semantic object labels and $3$D masks, unlike prior methods. In experiments, we demonstrate that FAST-Splat is $4$x to $6$x faster to train with a $13$x faster data pre-processing step, achieves between $18$x to $75$x faster rendering speeds, and requires about $3$x smaller GPU memory, compared to the best-competing semantic Gaussian Splatting methods. Further, FAST-Splat achieves relatively similar or better semantic segmentation performance compared to existing methods. After the review period, we will provide links to the project website and the codebase.
{shorinwa,jksun,schwager}@stanford.edu
1 Introduction
Research breakthroughs in vision-language foundation models have underpinned the remarkable performance of state-of-the-art object detection and classification [24, 50], segmentation [29, 15], and image captioning [27, 43]. In general, these models learn useful multi-modal image-language representations, entirely supervised with $2$D image-text pairs, all within a shared representation space. Recent work has shown that grounding the semantic knowledge encoded by vision-language foundation models in $3$D can be useful in improving semantic segmentation and localization [12, 35] (especially in models that take in low-resolution images as input), enabling applications such as $3$D scene-editing [14, 39] and robotic manipulation [28, 36]. Although these methods demonstrate compelling object segmentation performance from open-vocabulary queries, these methods fail to disambiguate between the semantic classes of unique objects that are semantically-similar to the natural language query [12]. For example, when a user queries for “tea,” existing semantic Gaussian Splatting methods localize the coffee machine without providing any clarifying information on the identity of the localized object, suggesting that “tea” exists in the scene, when the contrary is true. This ambiguity generally arises from the fact that many vision-language foundation methods, e.g., CLIP, utilize a bag-of-words approach in measuring the semantic similarity between image-language queries. Moreover, many semantic methods for Gaussian Splatting are notably slow, requiring a significant amount of memory and computation time for training and inference. We seek to derive a semantic Gaussian Splatting method that addresses these limitations.
In this work, we propose FAST-Splat, a Fast, Ambiguity-free Semantics Transfer method for Gaussian Splatting. FAST-Splat generates fine-grained semantic localization, with highly specific semantic classes even in the presence of ambiguous natural-language prompts, and as its name implies, runs much faster than existing semantic Gaussian Splatting methods. FAST-Splat utilizes the vision-language foundation model CLIP [27] to extend closed-set semantic Gaussian Splatting to an open-world setting, while retaining the non-ambiguity of localized objects inherent in closed-set methods. Unlike existing semantic Gaussian Splatting methods which rely on neural-field models for semantic distillation, FAST-Splat directly augments each Gaussian with specific semantic codes, without any neural model. This key insight not only enables FAST-Splat to exploit the explicit scene-representation of Gaussian Splatting for faster training and rendering speeds, but also enables FAST-Splat to provide unambiguous semantic object labels in response to open-vocabulary user queries, e.g., in semantic object localization, upon identifying semantically-similar objects using the semantic codes associated with each Gaussian, along with the semantic hash-table.
We compare FAST-Splat to existing semantic Gaussian Splatting methods [26, 49]. FAST-Splat achieves $4$x to $6$x faster training times and $18$x to $75$x faster rendering speeds, while requiring $3$x smaller GPU memory. Meanwhile, FAST-Splat achieves competitive semantic object localization performance, outperforming prior work in some scenes, while achieving the second-best performance in the other scenes. Notably, FAST-Splat enables semantic disambiguation in object localization, even with ambiguous user-provided natural-language queries, which we illustrate in LABEL:fig:teaser. In contrast to prior work, FAST-Splat provides a clarifying semantic label to each localized object, disambiguating the semantic identity of the localized object. For example, in LABEL:fig:teaser, when prompted with the query “tea,” FAST-Splat notifies a user that it localized a coffee machine, and an apple when given the query: “fruit.” Such disambiguation can be critical to enabling interesting downstream applications, e.g., in robotics.
To summarize our contributions:
  • We introduce FAST-Splat, a Fast, Ambiguity-Free Semantic Gaussian Splatting method that enables notably faster ($4$x to $6$x faster) language-semantics grounding in $3$D scenes, distilled from $2$D vision models.
  • FAST-Splat resolves semantic ambiguity in semantic object localization, arising from ambiguous user-provided natural-language queries or inherent scene ambiguity, enabling the identification of the precise semantic label of relevant objects.
  • FAST-Splat achieves superior rendering speeds for semantic Gaussian Splatting, about $18$x to $75$x faster rendering speeds compared to prior work.
2 Related Work
We provide a detailed review of existing work on open-vocabulary object detection and segmentation, along with follow-on work on grounding $2$D object detection and segmentation in $3$D by leveraging high-fidelity $3$D scene representations.
Open-Vocabulary Object Detection and Segmentation. The introduction of the vision transformer (ViT) architecture [6] has spurred rapid research advances in many classical computer vision tasks, such as object detection and segmentation. However, like their convolutional neural network (CNN) counterparts [31, 8], early ViT models [2, 45] were primarily trained to detect objects within a fixed number of classes, defined as closed-set object detection. Vision-image foundation models, e.g., CLIP [27] and transformer-based language encoders, such as BERT [5], have enabled the detection of objects through open-vocabulary natural-language prompts, without relying on a fixed number of object classes, referred to as open-set object detection. Open-vocabulary object detection methods [17, 22, 24, 44, 7] fuse the image and language embeddings from vision or language foundation models using a decoder transformer architecture with self-attention to detect objects in a query image. GLIP [17] and OWL-ViT [24] train an image and text encoder jointly to learn useful representations, which are fed to a decoder for $2$D grounding. While GLIP and OWL-ViT require a pretraining phase on large-scale internet image-text pairs, GroundingDINO [22] leverages pre-trained text encoder BERT [5], but otherwise utilizes a similar model architecture.
Zero-shot object segmentation methods [41, 21, 16, 29] leverage an encoder-decoder model architecture to generate a semantic mask of all instances of objects in an image. Mask DINO [16] introduces a mask prediction branch to DINO [45], extending it to object segmentation. Like GLIP, ViL-Seg [21] trains an image and text encoder together with a clustering head on image-caption pairs gathered on the internet to segment objects, enabling segmentation from natural-language prompts. SAM-2 [29] utilizes an image encoder with memory attention and a memory bank to enable video segmentation (in addition to image segmentation). However, SAM requires a mask, bounding box, or pixel-coordinate prompt for segmentation. GroundedSAM [32] extends SAM-2 to open-vocabulary queries by computing the bounding box associated with a given natural-language query, which is passed as an input prompt to SAM-2. Other methods [15, 18, 46] leverage pre-trained image and text encoders, which are fine-tuned in some cases, to enable open-vocabulary object segmentation.
Radiance Fields. Neural Radiance Fields (NeRFs) [23, 25, 1] represent a scene as a color and density field, parameterized by multi-layer perceptrons. NeRFs generate a photorealistic reconstruction of a scene, capturing high-fidelity visual details. However, high-quality NeRFs typically require long training times and achieve slow rendering rates. Gaussian Splatting [11] addresses these limitations by representing a scene using ellipsoidal primitives with visual attributes, such as opacity and color using spherical harmonics, in addition to the spatial and geometric attributes, such as the ellipsoid’s mean and covariance. Radiance fields can be trained entirely from RGB images, making them suitable for many vision tasks.
Grounding Language in 3D. To bridge the gap between $2$D object detection and $3$D object localization, prior work has examined grounding semantic embeddings from pre-trained image and text encoders in the $3$D world. CLIP-Fields [35] learns an implicit spatial mapping from $3$D points to CLIP image embeddings and Sentence-Bert [30] text embeddings using back-projected RGB-D points. VLMaps [10] and NLMap [3] distill vision-language features into grid-based maps. The sparsity of these maps limits the resolution of the open-vocabulary segmentation results achievable by these methods. For high-resolution object segmentation, Distilled Features Fields (DFF) [14], CLIP-NeRF [39], and LERF [12] ground CLIP [27], LSeg [15], and DINO [45] features in a NeRF representation of the scene. Moreover, similar techniques have been employed in [26, 9, 49, 51] to distill semantic features into Gaussian Splatting representations. The resulting distilled radiance fields (DRFs) enable dense open-vocabulary object segmentation in $3$D and have been applied to scene-editing [14, 39] and robotic manipulation [28, 36, 48, 37]. Despite their success in object segmentation, existing DRFs fail to precisely identify the semantic class of the segmented object, e.g., for a given query, these methods localize all objects related to the query without resolving the unique semantic identity of each object. This ambiguity creates a gap in scene understanding, which is often required in many downstream tasks, e.g., in robotic exploration or manipulation. In this work, we introduce a DRF with fine-grained semantics, enabling precise semantic identification of objects in a scene, disambiguating language-guided semantic object localization, and addressing existing challenges.
3 Preliminaries
Here, we provide some background information on Gaussian Splatting. Gaussian Splatting [11] represents a scene using ellipsoidal primitives, with ellipsoid $i$ parameterized by its mean ${\mu_{i}\in\mathbb{R}^{3}}$; covariance ${\Sigma=RSS^{\mathsf{T}}R^{\mathsf{T}}\in\mathbb{R}^{3\times 3}}$, where $R$ denotes a rotation matrix, and $S$ denotes a diagonal scaling matrix; opacity $\alpha$; and spherical harmonics (SH). Gaussian Splatting utilizes a tile-based rasterization procedure for efficient real-time rendering rates, outperforming NeRF-based representations, with the color $C$ of each pixel given by the $\alpha$-based blending procedure:
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}{i-1}(1-\alpha_{j}),$ (1)
where $c_{i}$ denotes the color of each ellipsoid and $\alpha_{i}$ denotes the per-ellipsoid opacity multiplied by the probability density of the $2$D Gaussian associated with the ellipsoid. Upon initialization from a sparse point cloud, which can be obtained from structure-from-motion, the attributes of each ellipsoid are optimized through gradient descent on a dataset of RGB images. Distilled feature fields build upon NeRFs and Gaussian Splatting, leveraging the same volumetric rendering and tile-based rasterization for NeRFs and Gaussian Splatting, respectively, to generate $2$D feature maps from the $3$D feature fields.
4 Fast, Ambiguity-Free Semantic Transfer
Now, we discuss FAST-Splat, a fast, fine-grained semantic distillation method for grounding $2$D language semantics in $3$D, enabling precise identification of the semantic label of objects in the $3$D scenes. We build off $3$D Gaussian Splatting to obtain a high-fidelity, photorealistic scene reconstruction entirely from RGB images. Given a dataset of images $\mathcal{D}$, Gaussian Splatting optimizes the mean, covariance, opacity, and SH parameters of each ellipsoid comprising the representation via the loss function:
$\mathcal{L}_{\mathrm{gs}}=(1-\lambda)\sum_{\mathcal{I}\in\mathcal{D}}\left\|\mathcal{I}-\hat{\mathcal{I}}\right\|_{1}+\lambda\mathcal{L}_{\mathrm{D-SSIM}},$ (2)
where ${\lambda\in[0,1]}$ denotes a weighting parameter between the ${\ell_{1}}$-loss term and the differentiable structural similarity index measure (D-SSIM) loss term, and $\mathcal{I}$ and $\hat{\mathcal{I}}$ denote the ground-truth and predicted RGB image, respectively. The Gaussian Splatting model is trained using the Adam optimizer [13], a gradient-based optimization method.
Almost all existing semantic distillation methods rely on neural architectures for grounding language in $3$D in NeRFs [39, 14, 12] and Gaussian Splatting scenes [26, 9, 49, 51]. Although effective in enabling open-set semantic object localization, neural-based approaches introduce significant computational bottlenecks into the training pipeline of NeRF and Gaussian Splatting scenes. Moreover, when utilized in Gaussian Splatting, neural-based semantic distillation essentially erases the computational gains, e.g., faster training times and rendering rates, afforded by Gaussian Splatting when compared to NeRFs.
Via FAST-Splat, we seek to derive a method for open-set semantic distillation, while preserving the faster training times and rendering rates achieved by the underlying Gaussian Splatting method. To achieve this goal, we revisit the fundamental insight exploited by the authors of the original Gaussian Splatting work [11], namely: that only representing occupied space explicitly can lead to significant speedup in training and rendering times without any notable degradation in the visual fidelity of the representation. Specifically, Gaussian Splatting harnesses the explicit form and sparsity of an ellipsoidal-primitive-based scene representation to achieve faster training times and rendering times compared to NeRFs, the state-of-the-art at the time, which relied on implicit fields for the color and density of the scene. By learning neural encoders for dimensionality reduction or neural fields for language semantics, existing semantic distillation methods for Gaussian Splatting fail to take advantage of this insight. In contrast, in FAST-Splat, we leverage this insight to speed up training and rendering times, by eschewing neural models. In particular, we augment the attributes of each ellipsoid with a semantic parameter, associated with its semantic identity ${\beta\in\mathbb{R}^{3}}$, which is trained simultaneously with the ellipsoid’s other attributes.
Closed-Set Semantic Gaussian Splatting. We formulate the open-set semantic distillation problem as the problem of extending closed-set semantic distillation (segmentation) methods to the open-set regime. To obtain the closed-set of semantic categories, we feed each image ${I\in\mathcal{D}}$ into an object detector, e.g., YOLO [40], DETR [2], and Efficient DETR [42], to generate a list of objects present in the image. Although closed-set object detectors generally exhibit remarkable performance, these detectors occasionally fail to identify all instances of an object in an image, given that the size of their dictionary of detectable objects is often limited. To alleviate this issue, we further process the outputs of the object detector using an open-set object detector, e.g., GroundingDINO [22], with the input prompt given by the classes of objects identified by the object detector. In preliminary results, we find that this additional data pre-processing stage improves the robustness of the data pre-processing pipeline, by eliminating false detections made by the closed-set object detector.
Figure 1: Unlike prior semantic Gaussian Splatting methods, FAST-Splat trains a non-neural-based semantic Gaussian Splatting model to achieve much faster training and rendering times by extending closed-set semantic segmentation to the open-vocabulary setting. FAST-Splat distills language semantics from a closed-set object detector, e.g., YOLO, augmented with an open-set object detector, e.g., GroundingDINO, and a image segmentation model, e.g., SAM-2.
At this stage, each image has a list of constituent objects and bounding boxes specifying the location of these objects in the image. However, high-fidelity semantic distillation generally requires pixel-wise semantic classes. To generate dense semantic maps, we utilize SAM-2 [29], passing in each object class and its associated bounding box. We maintain a dictionary (hash table), representing the closed-set of objects identified in the image. Now, we can distill the pixel-wise semantic information into the underlying Gaussian Splatting model. To predict the semantic class of each ellipsoid in the scene, we apply an affine transformation $\mathcal{W}$ to the semantic attribute $\beta$ of each ellipsoid, mapping $\beta$ to an $N$-dimensional space, where $N$ denotes the size of the dictionary. Subsequently, we apply the softmax function to the resulting outputs to define a probability over the entries in the dictionary. We extend this procedure to view-based rendering. Given a camera pose for rendering, we utilize differentiable tile-based rasterization [11] to render a semantic feature map corresponding to the camera pose. We utilize the aforementioned procedure to transform the semantic feature of each pixel to a probability distribution over the entries in the closed-set of object classes. We optimize the semantic attributes using the multi-class cross-entropy loss function, given by:
$\mathcal{L}_{\mathrm{ce}}=-\sum_{\mathcal{I}\in\mathcal{D}}\frac{1}{|\mathcal{T}|}\sum_{p\in\mathcal{T}}\gamma_{p}\log\left(\frac{\exp(\mathcal{F}_{\mathcal{I},p,c})}{\sum_{j=1}^{N}\exp(\mathcal{F}_{\mathcal{I},p,j})}\right),$ (3)
where $\mathcal{T}$ represents the set of all indices of pixels in the rendered feature map $\mathcal{F}_{\mathcal{I}}$ associated with image $\mathcal{I}$ (after applying the affine transformation $\mathcal{W}$), $c$ denotes the true class of the pixel, and $\gamma_{p}$ represents the weight associated with pixel $p$. We augment the set of object classes for each view with an undetected class label and assign a lower weight to pixels with this class label, i.e., pixels that were not segmented during the data pre-processing phase. This augmentation step serves as a regularization term for the stability of undetected pixels. We train all attributes of the Gaussian Splatting representation, e.g., the geometric, visual, and semantic attributes and the parameters in $\mathcal{W}$, simultaneously using the loss function: ${\mathcal{L}_{\mathrm{sgs}}=\mathcal{L}_{\mathrm{gs}}+\mathcal{L}_{\mathrm{ce}},}$ with $\mathcal{L}_{\mathrm{gs}}$ defined in (2).
Open-Vocabulary Semantic Gaussian Splatting. The resulting distillation method enables object localization from a pre-defined set of object classes. However, it does not support open-ended queries from users, which represents an important practical use-case. Here, we extend the capability of the distillation method from the closed-set setting to the open-vocabulary setting, leveraging open-source pre-trained text encoders, e.g., CLIP [27] and BERT [4]. These pre-trained text encoders compute text embeddings for natural-language inputs, mapping these inputs to a metric space.
We pre-compute the text embeddings $\theta_{\mathrm{d}}$ of each entry in the dictionary of detected object classes using the pre-trained text encoder, without any fine-tuning. At runtime, given a natural-language prompt specified by a user, FAST-Splat computes the text embedding $\theta_{\mathrm{q}}$ associated with the prompt using the pre-trained text encoder. To identify semantically-relevant objects in the scene, we compute the relevancy score [12], which represents the pairwise softmax between the cosine-similarity values computed over the dictionary embeddings $\theta_{\mathrm{d}}$ and a set of text embeddings consisting of the query embedding $\theta_{\mathrm{q}}$ (referred to as the positive embedding) and a set of canonical embeddings (referred to as negative embeddings).
Given a threshold on the relevancy score, we designate all entries whose relevancy score exceeds this specified threshold as being similar to the query. We map these entries to pixels in a rendered feature map using the probability distribution over all entries computed via the softmax function. In general, we can assign a label to each pixel by taking the maximum probability value; however, alternative techniques can be utilized, including those that provide a multi-hypothesis prediction set, e.g., conformal prediction [34].
Fine-grained Language Semantics. In the real-world, identifying only the location of a related object of interest, e.g., in semantic object localization, is generally inadequate. Identifying the precise semantic class of the localized object is of even greater importance, especially for downstream applications where semantic object localization constitutes a single component of a solution pipeline. Prior semantic Gaussian Splatting methods lack this critical capability. For example, existing semantic Gaussian Splatting methods will localize a coffee machine when prompted with the query “tea,” without disambiguating the semantic identity of the localized object, which could be misleading to a user. FAST-Splat seeks to address this limitation. The novel distillation technique employed by FAST-Splat endows the algorithm with the capability to not only identify the precise semantic class of each object, but also to disambiguate the semantic identity of related objects. Given a natural-language prompt, FAST-Splat identifies the related object classes from its dictionary and subsequently, provides the location of these objects in the scene along with their specific semantic label, illustrated in LABEL:fig:teaser. We summarize the architecture in Figure 1.
5 Experiments
We benchmark FAST-Splat against the primary, existing semantic Gaussian Splatting methods LangSplat [26] and Feature-3DGS [49]. Unfortunately, FMGS [51], another primary method, does not provide a publicly-available code implementation, preventing us from benchmarking against this method. We utilize the original code implementation provided by the authors of these baseline methods. In addition, we examine the capability of FAST-Splat in disambiguating the semantic identities of objects, e.g., in scenarios with ambiguous user prompts and scenarios with multiple related objects. We examine each method across five scenes: Kitchen; Cleaning; Meal; Workshop; and Plants. We abbreviate the name of each scene using the initial letter. We utilize YOLO [40] as the closed-set object detector; we provide additional implementation details and further discussion in the Supplementary Material.
Semantic Gaussian Splatting. We examine the fidelity of the $3$D scene reconstruction achieved by FAST-Splat in comparison to prior work LangSplat [26] and Feature-3DGS [49], beginning with a discussion of the computation time required for data pre-processing as well as the training time for each method, reported in Table 1. FAST-Splat requires at least a $13$x faster data pre-processing time compared to the best-competing method Feature-3DGS, and is about $78$x faster when compared to LangSplat. Similarly, FAST-Splat achieves about $5$x (and up to $8$x) faster training times compared to Feature-3DGS, and is about $4$x to $6$x faster to train compared to LangSplat.
Table 1: The data pre-processing time (prep.) and training time (train) for the semantic Gaussian Splatting methods across five scenes. FAST-Splat outperforms all existing methods.
Scene Method Prep. (min) $\downarrow$ Train (min) $\downarrow$ K. LangSplat [26] 128.49 58.54 Feature-3DGS [49] 15.09 73.15 FAST-Splat [ours] 1.12 12.08 C. LangSplat [26] 85.53 73.32 Feature-3DGS [49] 16.49 85.10 FAST-Splat [ours] 1.01 12.44 M. LangSplat [26] 82.67 68.22 Feature-3DGS [49] 16.36 68.01 FAST-Splat [ours] 1.06 11.71 W. LangSplat [26] 103.20 66.32 Feature-3DGS [49] 15.17 65.27 FAST-Splat [ours] 0.81 12.18 P. LangSplat [26] 136.99 62.61 Feature-3DGS [49] 16.05 110.86 FAST-Splat [ours] 1.23 13.79
We utilize standard metrics for comparing $3$D reconstruction methods, such as the peak-signal-to-noise ratio (PSNR) and Structural Similarity Index Measure (SSIM) for evaluating the reconstruction quality; the training time; the memory usage; and the rendering frame-rate. In Table 2, we report the performance of each method on these metrics. FAST-Splat outperforms all other methods on the rendering frame-rate and memory-usage metrics, achieving the highest rendering speeds while requiring the least amount of GPU memory. In four of the five scenes, FAST-Splat achieves the highest SSIM score, outperforming the next best-competing method LangSplat. LangSplat achieves the highest PSNR score in four of the five scenes, followed by FAST-Splat. FAST-Splat achieves the highest PSNR score in the Plants scene. Further, FAST-Splat reduces the memory usage in semantic Gaussian Splatting by a factor of more than $3$x compared to the best-competing method LangSplat and a factor of more than $6$x compared to Feature-3DGS. In essence, FAST-Splat achieves a higher-fidelity scene reconstruction with a lower memory usage.
Lastly, we present the rendering speed of each method. FAST-Splat attains $18$x to $75$x faster rendering speeds compared to the best-competing method LangSplat. Similarly, FAST-Splat achieves about $550$x faster-rendering frame-rates compared to Feature-3DGS, highlighting the superior performance of FAST-Splat. We show rendered images from each of the scenes with the different methods in Figure 2, juxtaposed with the ground-truth RGB image. We render novel views in FAST-Splat.
Table 2: The rendering frame-rate (Speed), image quality, and memory usage (Memory) achieved by the semantic Gaussian Splatting methods across the five scenes.
Scene Method Speed $\uparrow$ (F/s) SSIM $\uparrow$ PSNR $\uparrow$ Memory $\downarrow$ (GB) K. LangSplat [26] 2.25 0.966 35.26 8.05 Feature-3DGS [49] 0.23 0.931 30.20 7.72 FAST-Splat [ours] 115.23 0.971 34.24 1.02 C. LangSplat [26] 2.25 0.918 32.51 9.54 Feature-3DGS [49] 0.22 0.871 28.70 7.72 FAST-Splat [ours] 114.37 0.929 30.35 1.00 M. LangSplat [26] 2.24 0.933 33.37 9.54 Feature-3DGS [49] 0.22 0.845 27.83 7.86 FAST-Splat [ours] 127.41 0.940 31.25 0.98 W. LangSplat [26] 1.87 0.921 31.41 10.08 Feature-3DGS [49] 0.22 0.357 8.11 7.72 FAST-Splat [ours] 141.48 0.917 29.22 1.04 P. LangSplat [26] 6.84 0.915 28.79 7.09 Feature-3DGS [49] 0.21 0.773 25.33 7.72 FAST-Splat [ours] 128.77 0.942 31.12 1.17
Figure 2: RGB images rendered using each semantic Gaussian Splatting method. FAST-Splat achieves $18$x to $75$x faster rendering speeds with at least $3$x lower memory usage, while achieving competitive (and in some cases, better) reconstruction quality.
Semantic Object Localization. Now, we evaluate each method on semantic object segmentation from open-vocabulary queries, using the mean intersection-over-union (mIoU) and accuracy metrics. In Table 3, we report the performance of each method on these metrics. FAST-Splat achieves the best performance in two scenes: the Kitchen scene and the Workshop scene. In the Plants scene, Feature-3DGS achieves the highest mIoU score, while LangSplat achieves the highest accuracy. Further, in the Cleaning scene, Feature-3DGS achieves the highest mIoU score, while FAST-Splat achieves the highest accuracy. In all scenes where FAST-Splat does not achieve the best score, FAST-Splat achieves the next-best score, remaining competitive with the best-performing method, which underscores the overall effectiveness of our proposed method. In Figure 3, we show segmentation masks for each of the methods across the five scenes.
Notably, all methods perform poorly in the Workshop scene. We believe this observation may be due to the Workshop scene being mostly out-of-distribution, considering that the training-data distribution of most vision models and vision-language models do not contain industrial objects, such as those found in workshops. Recall that we utilize the YOLO model for closed-set object detection. The YOLO model is trained using the object categories found in the COCO dataset [19], which does not contain common objects found in a workshop, such as a power drill and screwdriver, which is present in the Workshop scene. Moreover, other open-vocabulary object detection methods also struggle to identify industrial objects, e.g., GroundingDINO [22].
Figure 3: Semantic segmentation using each semantic Gaussian Splatting method. FAST-Splat achieves competitive performance compared to existing semantic Gaussian Splatting method, achieving the highest mIoU and accuracy scores in some scenes while being the next best-performing method in the remaining scenes.
Table 3: The mean intersection-over-union (mIoU) and accuracy of the segmentation masks generated by semantic Gaussian Splatting methods across the five scenes.
Scene Method mIoU $\uparrow$ Accuracy $\uparrow$ K. LangSplat [26] 0.525 0.920 Feature-3DGS [49] 0.550 0.920 FAST-Splat [ours] 0.709 0.925 C. LangSplat [26] 0.552 0.850 Feature-3DGS [49] 0.570 0.751 FAST-Splat [ours] 0.517 0.901 M. LangSplat [26] 0.409 0.714 Feature-3DGS [49] 0.571 0.913 FAST-Splat [ours] 0.525 0.904 W. LangSplat [26] 0.054 0.350 Feature-3DGS [49] 0.061 0.327 FAST-Splat [ours] 0.127 0.908 P. LangSplat [26] 0.562 0.900 Feature-3DGS [49] 0.641 0.872 FAST-Splat [ours] 0.570 0.861
Figure 4: FAST-Splat resolves language ambiguity in user-provided natural-language queries in semantic object localization. FAST-Splat identifies the specific semantic class of each relevant object, e.g., a coffee machine and a kettle, when prompted with an ambiguous query, e.g., “coffee” and “cooking pot,” respectively. Likewise, FAST-Splat disambiguates between a “cup” and a vase and between a “fruit” and a pottedplant in the garden-like scene.
Semantic Disambiguation. FAST-Splat extends the capabilities of semantic Gaussian Splatting methods beyond semantic object localization to the domain of semantic disambiguation. As noted in the preceding discussion, existing semantic Gaussian Splatting methods cannot resolve semantic ambiguity either arising from the user-provided natural-language query or inherently from the scene composition. In contrast, when provided with an ambiguous natural-language query, FAST-Splat disambiguates the semantic object localization task, by identifying the specific semantic label of related objects in the scene.
We provide some examples in Figure 4. For example, when a user queries for “coffee” in the Kitchen scene, FAST-Splat provides the user with the semantic identity of the related object coffee machine, along with the semantic segmentation mask and location of the coffee machine, given a threshold for the segmentation task. Likewise, when queried for a “cooking pot,” FAST-Splat identifies the semantic class: kettle, and the segmentation mask and location for the kettle. Notably, in the Plants scene, FAST-Splat resolves the ambiguity between a cup and a vase. When queried with the prompt “cup,” FAST-Splat informs the user that, more precisely, a vase exists in the scene. Similarly, when asked to localize a “fruit,” FAST-Splat clarifies that the scene contains a pottedplant and provides the semantic segmentation mask and location.
Scene Editing. Further, like prior semantic Gaussian Splatting methods, FAST-Splat enables scene-editing of Gaussian Splatting scenes. With FAST-Splat, a user can modify the visual properties of objects in the scene, and in addition, can insert, delete, or move objects in the scene. For space considerations, we omit a detailed discussion of this capability, given that it has been discussed in prior work. In Figure 5, we provide an example where the visual properties of the coffee machine is modified.
Figure 5: Like existing semantic Gaussian Splatting methods, FAST-Splat enables editing of Gaussian Splatting scenes. Here, we show color-editing of a coffee machine in a kitchen.
6 Conclusion
We introduce FAST-Splat, a fast semantic Gaussian Splatting method that resolves ambiguity in semantic object localization. FAST-Splat retains the explicit form of Gaussian Splatting and avoids neural architectures to achieve fast training and rendering speeds compared to existing methods. Further, FAST-Splat builds upon closed-set semantic distillation to enable open-vocabulary semantic Gaussian Splatting. We leverage this formulation to endow FAST-Splat with the capability to resolve semantic ambiguity arising from the natural-language query or the scene composition, in the semantic object localization task, enabling FAST-Splat to address one of the fundamental limitations of existing semantic Gaussian Splatting methods.
7 Limitations and Future Work
As with existing semantic Gaussian Splatting methods, the performance of FAST-Splat is highly impacted by the detection performance of the closed-set object detector. Currently, the size of the object categories detected by open-source closed-set object detectors is quite limited. For example, YOLO [40] can only detect $80$ objects, which is quite small, considering the diversity of objects in the real-world. This limitation is not imposed by the underlying model architecture. In fact, YOLO can be trained on a much greater diversity of objects. However, the model was trained using the COCO dataset, introducing this limitation. In comparison, existing vision-language models are trained on more diverse datasets, enabling these models to detect a much broader range of objects. Future work will seek to utilize closed-set object detectors trained with a much larger set of object categories.
Our work utilizes an open-vocabulary object detector, e.g., GroundingDINO [22], in addition to a closed-set object detector, which enables FAST-Splat to mitigate the limitations of the closed-set object detector. Essentially, we can always augment the set of detectable objects through natural-language when querying the open-vocabulary object detector. In fact, we are currently exploring using image-tagging models, e.g., RAM [47], or multi-modal vision-language models, e.g., LLaVa [20] to detect objects for language grounding. Future work will seek to explore this research direction.
Acknowledgment
This work was supported in part by NSF project FRR 2342246, and ONR project N00014-23-1-2354, and DARPA project HR001120C0107. Toyota Research Institute provided funds to support this work.
Appendix A Implementation Details
To implement FAST-Splat, we build off Splatfacto in Nerfstudio [38], and utilize the text encoder of the OpenAI’s CLIP ResNet model RN50x64 [27] to compute the text embeddings. However, other text encoders can also be used, e.g., other CLIP models or BERT [4]. We train FAST-Splat using Stochastic Gradient Descent with momentum, via the Adam optimizer [13].
We collected the training data by taking videos of the scene using a smartphone camera. We utilized structure-from-motion, e.g., COLMAP [33], for computing the pose of the camera for each frame used in training the Gaussian Splatting model. FAST-Splat is trained on an NVIDIA GeForce RTX 3090 GPU with 24GB VRAM for $30000$ iterations, using the hyperparameters specified in [38] in its Splafacto model. We train the baseline methods LangSplat and Feature-3DGS using the original code implementation provided by their respective authors on an NVIDIA GeForce RTX 3090 Ti GPU with 24GB VRAM, giving LangSplat and Feature-3DGS an advantage on hardware, since the 3090 Ti has $10752$ CUDA cores compared to the $10496$ CUDA cores on the 3090, among other improved hardware specifications. To train LangSplat, we follow the procedure provided by its authors, which involves pre-training the Gaussian Splatting Scene using the code implementation provided in [11], prior to training the semantic components. We do not include the structure-from-motion processing times in any of the results reported in the experiments, given that this procedure is required universally by all methods.
Appendix B Semantic Object Localization
Here, we provide additional results, demonstrating the photorealistic novel-view synthesis capability of FAST-Splat, including its semantic segmentation masks. In Figure 6, we show the semantic localization of a “wall,” “countertop,” “kettle,” and “cooking spoon.” In each case, FAST-Splat successfully localizes the pertinent object. The colors in each figure vary with the similarity of the object to the natural-language query. We utilize the “Turbo” colormap, with an increase in the intensity of the red color channel indicating an increase in the similarity of an object to the query and an increase in the intensity of the blue color channel indicating a decrease in the relevance of an object to the query, in general. Moreover, the RGB images rendered by FAST-Splat appear photorealistic, displayed in the top row in Figure 6, all rendered at novel views.
Figure 6: FAST-Splat enables semantic object localization wth open-vocabulary user prompts. Here, we provide segmentation results for a “wall,” “countertop,” “kettle,” and “cooking spoon.”
Appendix C Semantic Disambiguation
As discussed earlier in this work, FAST-Splat enables semantic disambiguation, e.g., in semantic object localization. Not only does FAST-Splat provide the semantic identity of the most relevant object, FAST-Splat also identifies other pertinent objects, providing the semantic label of these objects along with their similarity to the natural-language prompt. In Figure 7, we demonstrate the semantic disambiguation capabilities of FAST-Splat. In the top row, when queried for a “fork,” FAST-Splat identifies the spoon as the object that is most similar to the query. In addition, FAST-Splat localizes the knife, identifying it as being relevant, although less similar, compared to the spoon. Likewise, in the bottom row, when prompted with the ambiguous query “water,” FAST-Splat localizes the kettle, noting its lower similarity compared to a coffee machine, which FAST-Splat also localizes in the scene. Lastly, FAST-Splat identifies the bottle as the most similar object to the query, resolving the ambiguous prompt.
Figure 7: FAST-Splat resolves semantic ambiguity in object localization from natural-language queries. In the top row, although the queried object “fork” does not exist in the scene, FAST-Splat localizes relevant objects in the scene, and more importantly, provides the semantic label of each of these objects. Here, FAST-Splat identifies a spoon and a knife and notes that the spoon (not the knife) is more similar to the fork. Likewise, in the bottom row, given the ambiguous prompt “water,” FAST-Splat localizes a kettle, coffee machine, and bottle, providing their semantic object classes along with their relative similarity to the prompt.
Appendix D Limitations
We discuss the limitations of FAST-Splat further. The performance of FAST-Splat may degrade in scenes that are out of the distribution of the training data for the closed-set or open-set object detector used in extracting the object classes. For example, the workshop scene lies primarily outside the distribution of the training data of YOLO [40]. Consequently, FAST-Splat performs relatively poorly in this scene. Figure 8 shows the semantic localization results for a “spatula” and “tape.” FAST-Splat fails to localize these objects accurately. However, this limitation can be mitigated by training the closed-set detector on a broader set of object classes; augmenting the query passed into the open-vocabulary object detector; or training the open-vocabulary object data on more data, depending on the primary cause.
Figure 8: The performance of FAST-Splat degrades in scenes that are out-of-distribution for the object detector. Here, FAST-Splat fails to generate accurate segmentation masks in the Workshop scene, which contains objects that are not found in the training dataset of many object detectors.